# Emergence of KNO scaling in multiplicity distributions in jets produced at the LHC

### v1

@@ I Introduction
One of the remarkable features exhibited by MDs is the famous Koba-Nielsen-Olsen (KNO) scaling kno ; kno-1 ; kno-2 ; kno-3 , a phenomenon expected to be observed at asymptotically high energies. This prediction was made before the existence of QCD. Later there were several attempts to understand it in terms of quark and gluon dynamics, such as in Refs. dumi12 ; dumi13 . In these works it was shown that KNO scaling emerges if the effective theory describing color charge fluctuations at a scale on the order of the saturation momentum is approximately Gaussian. Moreover both non-linear saturation effects and running-coupling evolution are required in order to obtain KNO scaling. Very recently, in Ref. kno1 a MD satisfying KNO scaling was derived by solving the Mueller dipole evolution equation in the double logarithm approximation. This supports the idea that gluon emission is a Markov process in which the emitted gluons are strongly ordered in rapidity.
@@ II Revisiting the ATLAS data
Figure 2: Variance as a function of a) $<n>$ and b) $p_{T}$. Data are from Refs. atlas11 and atlas19 . The solid line shows the result obtained with the Poisson distribution.
@@
In Fig. 1 we plot the ATLAS data on jet multiplicities in the KNO form. In Fig. 1a and Fig. 1b we show the data from Ref. atlas11 on lower $p_{T}$ jets. The two sets refer to two values of the jet $R$ variable. We clearly see that the data violate KNO scaling. However, if we consider the higher $p_{T}$ jets measured in Ref. atlas19 we observe the onset of scaling, as shown in Fig. 1c, specially at $p_{T}>300$ GeV. This change of behavior can be more clearly seen if we plot the ratio between the variance and the average multiplicity ($Var(n)/<n>$) which is equal to one for a Poisson distribution. In Fig. 2 we can see that this ratio is below one for lower $<n>$ (Fig. 2a) and for lower $p_{T}$ (Fig. 2b) and around $<n>\simeq 10$ or $p_{T}\simeq 30$ GeV there is a clear change. The ratio becomes larger than one and we go from sub-poissonian to a super-poissonian distribution. It is tempting to associate this broadening of the multiplicity distribution with transition from quark to gluon initiated jets. In Ref. opal98 the properties of quark and gluon jets were studied. In particular, it was found that the dispersion D of the multiplicity distribution from jets was $D_{g}\simeq 4.37$ and $D_{q}\simeq 4.30$ for gluon and quark jets respectively. The errors quoted in opal98 are very large but these numbers suggest that gluon jets are broader than quark ones and the onset of the dominance of the former could be the dynamical cause of the behavior observed in Fig. 2.
@@
$<n_{ch}>=a\,[\alpha_{s}(p_{T})]^{b}\,e^{c/\sqrt{\alpha_{s}(p_{T})}}\,[1+d\,\sqrt{\alpha_{s}(p_{T})}]$ (1)
@@
$<n_{ch}>=\mathcal{K}_{ch}\,exp\,\left[2.50217\,\sqrt{Y}-0.491546\,ln\,Y-(0.06889-0.41151\,ln\,Y)\frac{1}{\sqrt{Y}}+(0.00068-0.161658\,ln\,Y)\frac{1}{Y}\right]\\ \$ (2)
@@ III The Sub-Poissonian distribution
$\lambda_{n}=\lambda{(n+1)^{-\delta}}+\lambda_{0}\quad{\rm and}\quad\mu_{n}=n\mu,$ (5)
@@
$-\left[\lambda(n+1)^{-\delta}+\lambda_{0}+n\mu\right]P(n)+(\lambda n^{-\delta}+\lambda_{0})P(n-1)+(n+1)\mu P(n+1)=0$ (6)
@@
$\frac{\lambda}{\mu}=\alpha\quad{\rm and}\quad\frac{\lambda_{0}}{\mu}=\alpha_{0},$ (7)
@@ IV From Sub-Poisson to Negative Binomial Distribution
The fits are very good. They show that $k$ decreases with the jet $p_{T}$ in the same way as it decreases with the energy in NBD fits of the minimum bias multiplicity distributions ge22 . They also show that $\bar{n}/k$ increases with $p_{T}$ indicating the approach to KNO scaling, which is reached when $\bar{n}>>k$. In fact, it was shown in dumi12 that when this ratio reaches 66~{}66 one already has a very good scaling. From the last entries of Table 5 we have ratios close to this number.
@@
A detailed interpretation of these results in terms of the QCD dynamics is beyond the scope of this work but there are hints which may help theorists. The decreasing values of $\alpha$ suggest that the death rate must increase with the energy. In a parton cascade this would mean that the process $gg\to g$ becomes more important and when this happens we are approaching the gluon saturation regime. The initially positive values of $\delta$ render $P(n)$ narrow. This may be a consequence of phase space restrictions. At lower energies we have a smaller number of particles and energy-momentum conservation prevents large fluctuations. At higher energies and larger number of particles the fluctuations are also larger and $P(n)$ becomes broader. Alternatively, the narrowness of $P(n)$ at lower values of $p_{T}$ may indicate that the jets are initiated by quarks. A broader $P(n)$ would indicate the dominance of gluon initiated jets. Further investigation is required.

### v2

@@ I Introduction
One of the remarkable features exhibited by MDs is the famous Koba-Nielsen-Olsen (KNO) scaling kno ; kno-1 ; kno-2 ; kno-3 , a phenomenon expected to be observed at asymptotically high energies. This prediction was made before the existence of QCD. Later there were several attempts to understand it in terms of quark and gluon dynamics, such as in Refs. dumi12 ; dumi13 . In these works it was shown that KNO scaling emerges if the effective theory describing color charge fluctuations at a scale of the order of the saturation momentum is approximately Gaussian. Moreover both non-linear saturation effects and running-coupling evolution are required in order to obtain KNO scaling. Very recently, in Ref. kno1 a MD satisfying KNO scaling was derived by solving the Mueller dipole evolution equation in the double logarithm approximation. This supports the idea that gluon emission is a Markov process in which the emitted gluons are strongly ordered in rapidity.
@@ II Revisiting the ATLAS data
Figure 2: Variance as a function of a) $\bar{n}$ and b) $p_{T}$. Data are from Refs. atlas11 and atlas19 . The solid line shows the result obtained with the Poisson distribution.
@@
In Fig. 1 we plot the ATLAS data on jet multiplicities in the KNO form. In Fig. 1a and Fig. 1b we show the data from Ref. atlas11 on lower $p_{T}$ jets. The two sets refer to two values of the jet $R$ variable. We clearly see that the data violate KNO scaling. However, if we consider the higher $p_{T}$ jets measured in Ref. atlas19 we observe the onset of scaling, as shown in Fig. 1c, specially at $p_{T}>300$ GeV. This change of behavior can be more clearly seen if we plot the ratio between the variance and the average multiplicity ($Var(n)/\bar{n}$) which is equal to one for a Poisson distribution. In Fig. 2 we can see that this ratio is below one for lower $\bar{n}$ (Fig. 2a) and for lower $p_{T}$ (Fig. 2b) and around $\bar{n}\simeq 10$ or $p_{T}\simeq 30$ GeV there is a clear change. The ratio becomes larger than one and we go from a sub-poissonian to a super-poissonian distribution. It is tempting to associate this broadening of the multiplicity distribution with the transition from quark to gluon initiated jets. In Ref. opal98 the properties of quark and gluon jets were studied. In particular, it was found that the dispersion D of the multiplicity distribution from jets was $D_{g}\simeq 4.37$ and $D_{q}\simeq 4.30$ for gluon and quark jets respectively. The errors quoted in opal98 are very large but these numbers suggest that gluon jets are broader than quark ones and the onset of the dominance of the former could be the dynamical cause of the behavior observed in Fig. 2.
@@
$\bar{n}_{ch}=a\,[\alpha_{s}(p_{T})]^{b}\,e^{c/\sqrt{\alpha_{s}(p_{T})}}\,[1+d\,\sqrt{\alpha_{s}(p_{T})}]$ (1)
@@
$\bar{n}_{ch}=\mathcal{K}_{ch}\,exp\,\left[2.50217\,\sqrt{Y}-0.491546\,ln\,Y-(0.06889-0.41151\,ln\,Y)\frac{1}{\sqrt{Y}}+(0.00068-0.161658\,ln\,Y)\frac{1}{Y}\right]\\ \$ (2)
@@ III The Sub-Poissonian distribution
$\lambda_{n}=\lambda{(n+1)^{-\delta}}+\sigma\quad{\rm and}\quad\mu_{n}=n\mu,$ (5)
@@
$-\left[\lambda(n+1)^{-\delta}+\sigma+n\mu\right]P(n)+(\lambda n^{-\delta}+\sigma)P(n-1)+(n+1)\mu P(n+1)=0$ (6)
@@
$\frac{\lambda}{\mu}=\alpha\quad{\rm and}\quad\frac{\sigma}{\mu}=\alpha_{0},$ (7)
@@ IV From Sub-Poisson to Negative Binomial Distribution
The fits are very good. They show that $k$ decreases with the jet $p_{T}$ in the same way as it decreases with the energy in NBD fits of the minimum bias multiplicity distributions ge22 . They also show that $\bar{n}/k$ increases with $p_{T}$ indicating the approach to KNO scaling, which is reached when $\bar{n}>>k$. Indeed, for KNO scaling, we expect the moments of the $P(z=n/\bar{n})$ distribution to be independent of $\bar{n}$. In the case of the NBD distribution, the second central moment of the $P(z)$ distribution is
$Var(z)=\frac{Var(n)}{\bar{n}^{2}}=\frac{1}{\bar{n}}+\frac{1}{k},$ (17)
For $\bar{n}\gg k$ we have an approximate scaling. In fact, it was shown in dumi12 that when this ratio reaches 66~{}66 one already has a very good scaling. From the last entries of Table 5 we have ratios close to this number.
@@
A detailed interpretation of these results in terms of the QCD dynamics is beyond the scope of this work but there are hints which may help theorists. We now present a few heuristic remarks, but further investigation is required for definite conclusions. They are the following:
$(i)$ Parton Saturation
The decreasing values of $\alpha$ suggest that the death rate must increase with the energy. In a parton cascade this would mean that the process $gg\to g$ becomes more important and when this happens we are approaching the gluon saturation regime. From Fig. 3 we see that the deviations from the standard perturbative QCD calculations occur at 11~{}11 TeV and this energy is high enough for saturation effects to become visible.
$(ii)$ Quark and gluon jets and phase space
The initially positive values of $\delta$ render $P(n)$ narrow. This may be a consequence of phase space restrictions. At lower energies we have a smaller number of particles and energy-momentum conservation prevents large fluctuations. At higher energies and larger number of particles the fluctuations are also larger and $P(n)$ becomes broader. Alternatively, the narrowness of $P(n)$ at lower values of $p_{T}$ may indicate that the jets are initiated by quarks. A broader $P(n)$ would indicate the dominance of gluon initiated jets.
$(iii)$ Threshold effects
Based on the ”parton liberation” picture AHM and on the ”local parton-hadron duality” DKTM we expect the $P(n)$ obtained from the parton cascade to be similar to the $P(n)$ of observed hadrons. In the parton model the average number of partons is related to the deep inelastic structure function and at high energies it is assumed that $\bar{n}\,\propto\,G(x,Q^{2})$, where $G$ is the gluon distribution. An increase in the number of gluons implies that the total momentum will be partitioned among more gluons and it becomes more likely to find a gluon with a small fraction of the total momentum. Particle production in the fusion of two gluons with momenta $x_{1}$ and $x_{2}$ introduces a threshold $x_{1}\,x_{2}\geq m^{2}/s$, where $m$ is the (sum of the) mass(es) of the produced hadron(s) and $\sqrt{s}$ is the proton-proton collision energy. Imposing this restriction will exclude the lower $x$ domain of $G(x)$ where the number of gluons grows rapidly and thus will exclude the larger $n$ configurations. In other words, the imposition of particle production in a collision changes the solution of the parton cascade equation, favoring a narrower $P(n)$ with smaller $\bar{n}$. This motivates the use of the power index $\delta$ in the birth rate in Eqs. (5) and in (14). As the energy $\sqrt{s}$ increases the threshold constraint becomes less restrictive and allows for a larger number of gluons producing a larger number of final hadrons. As a consequence $P(n)$ becomes broader and with a larger $\bar{n}$. In Eq. (14) this behavior translates into a decreasing value of $\delta$.
$(iv)$ Convolution, substructures and scaling
In minimum bias proton-proton collisions typically half of the energy is released in the central rapidity region and the other half is carried by the remnants of the incoming protons. The fraction of the energy released in the central region may change with the energy igm . This picture would suggest that the observed hadrons in the final state come from three sources. However the number of sources may be larger because the ”central fireball” may be composed of subsystems, smaller fireballs. This depends on details of the dynamics, i.e., perturbative or non-perturbative, with more or less string formation and decay, with or without thermalization, etc. In any case an important part of our understanding of the collision is to characterize the sources.
The number of sources will follow a distribution $F$ and each source will emit a number of hadrons with a distribution $G$. Therefore the final multiplicity distribution $P$ will be the result of the convolution $P=F\otimes G$. Alternatively, we can fix the number of sources to be one (which might be appropriate for jets) and let the average number of hadrons produced from this source, $\bar{n}$, to fluctuate according to a distribution $F$. Along this line, in Ref. PB (see also SM ) it was shown that, P will follow KNO scaling (at large n and for finite $n/\bar{n}$) if it can be written as a convolution where G is a Poisson distribution and $F$ a gamma function. In this case, the convolution $F\otimes G$ yielded a negative binomial distribution.
At lower energies, we do not observe scaling. On the other hand, at higher jet energies we observe empirically that KNO scaling is reached and also that the negative binomial distribution reproduces the data very well. This suggests that the limiting $P(n)$ for jets can be written as the convolution used in PB , allowing us to make conjectures about the fluctuations of $\bar{n}$ and about the nature of the sources.